OpenStudy (anonymous):

Explain, in complete sentences, how you would completely factor 15x2 – 25x – 60 by using the grouping method. How would you check your factors for accuracy?

OpenStudy (anonymous):

I am hopelessly lost when it comes to factor grouping.

OpenStudy (anonymous):

I just read a little bit about it, it still sounds like algeblah.

OpenStudy (anonymous):

not sure what you mean by grouping first thing here is to divide the function by 5: 5 (3x^3 - 5x - 12) now you need to find two numbers whose product is -12 and where the result of multiply by 3 and adding give the middle number -5 i do this by trial and error - 3 * 4 = -12 and -3*3 + 4 = -5 so the factors of the expression in the brackets is (3x + 4)(x - 3) ( + 4x - 9x = -5x) and final result is 5(3x + 4)(x - 3)

OpenStudy (anonymous):

I don't understand what they mean by grouping either. This is the site I was told to read - http://www.mathwarehouse.com/algebra/polynomial/how-to-factor-by-grouping.php

OpenStudy (anonymous):

ok - its similar way to mine - i suppose i've got into a groove with these and cant get out of it! i believe the best way to grasp it is to practice as many as you can the expressions starting with x^2 are easier than those with ar^2 eg factor x^2 - 2x - 15 - here we need 2 numbers whose product is - 15 and who add up to -2 these are -5 and 3 because -5 * 3 = -15 and -5 + 3 = -2 so factors are (x - 5)(x + 3)

OpenStudy (anonymous):

So it's basically the same thing?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

Oh crap, maybe I learned the old way wrong and I was learning this the whole time.

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Thanks man.

OpenStudy (anonymous):

yw - good luck